OpenStudy (anonymous):

Determine whether y varies directly with x. If so find the constant of variation k and write the equation. x y 7 11 8 13 9 15 10 17 Can someone please help me and explain

jimthompson5910 (jim_thompson5910):

If y varied directly with x, then y = k*x for some fixed constant number k

jimthompson5910 (jim_thompson5910):

Using this info, what is k?

OpenStudy (anonymous):

I don't know. How do you find it? I really don't understand this.

jimthompson5910 (jim_thompson5910):

The idea is that if you multiply each x term by some fixed number (the same fixed number) and you get the corresponding y value, then this is a direct variation Here's a simple example x y ------ 1 2 2 4 3 6 If we multiply each x by 2 we'll get the correct corresponding y value, so this is a direct variation

jimthompson5910 (jim_thompson5910):

y = k*x 11 = k*7 Now solve for k

OpenStudy (anonymous):

Oh that makes more sense! Okay (:

jimthompson5910 (jim_thompson5910):

So what is k when you solve that last equation?

OpenStudy (anonymous):

Could I do 11/7? K=1.57

jimthompson5910 (jim_thompson5910):

So if this was a direct variation, then we should be able to multiply each x by 11/7 or roughly 1.57 to get the corresponding y value But... When x = 8, k*x is (11/7)*8 = 88/7 which is roughly 12.57 and NOT 13 So this can't be a direct variation

OpenStudy (anonymous):

Since it's not, what am I supposed to do? Just leave it like that or..

jimthompson5910 (jim_thompson5910):

Just say it's not a direct variation. If it was a direct variation, then you would proceed to find k or show/say what k is.

OpenStudy (anonymous):

Oh alright

jimthompson5910 (jim_thompson5910):

since it says "If so find the constant of variation k and write the equation"

OpenStudy (anonymous):

jim is right and he proved that there is no direct constant for proportionality between y and x, so you don't have to do anything else but state that. There IS an equation (actually lots of them) that you could write to show a relationship between y and x, but you are not required to find one since none will exhibit a directly proportional relationship between y and x.